OpenStudy (anonymous):

Find the integral. 2 / x^2

OpenStudy (calculusfunctions):

Just giving away the answers? Which equals not helping? Even that much should not have been said.

OpenStudy (anonymous):

\[\frac{ 2 }{ x^2 }\]

OpenStudy (calculusfunctions):

\[\int\limits_{}^{}\frac{ 2 }{ x ^{2} }dx\]Correct?

OpenStudy (anonymous):

Ah yes, that what i was trying to type. Thanks.

OpenStudy (calculusfunctions):

This is a real easy one. Do have any idea of what would be the most efficient method here?

OpenStudy (anonymous):

Could I move the X^2 up and make it X^-2 ?

OpenStudy (calculusfunctions):

Yes! So please show me that initial step.

OpenStudy (anonymous):

Then add + 1 to the exponent and -2 + 1 to the denominator ?

OpenStudy (calculusfunctions):

Yes! You got it. Finish and show me the complete proper solution please.

OpenStudy (anonymous):